OpenStudy (anonymous):

Three 2.50 µC charges are placed on tiny conducting spheres at the ends of 1.00 m-long strings that are connected at 120º angles as shown below. The magnitude, in N, of the force on any one of the charges is

OpenStudy (irishboy123):

draw it?

OpenStudy (anonymous):

 sorry i cant draw nicely. it is supposed to be 3 straight line with circles on each end

OpenStudy (anonymous):

hope you can help

OpenStudy (irishboy123):

 the 3 charges should form an equilateral triangle of side \(\sqrt{3}\). there is also quite some symmetry in the set up. the force between the charge at the top and each of the other two charges will follow Coulumb's Law. For each pair, \(\large F = k \dfrac{q^2}{r^2}\) where \(\large q = 2.5 \mu F, r = \sqrt{3} \) in the left-right direction, these force will cancel -- due to symmetry. in the up-down direction we can add their **up-down components** so the overall force is \(\large F_{net} = 2 \times k \dfrac{q^2}{r^2} \cdot \sin 60^o\)